# Cooperative Bistatic ISAC Systems for Low-Altitude Economy

### v1

@@ II-B Sensing Channel Model
where $\alpha_{k}$ represents the composite channel gain of the $k$th path, while ${\delta}^{\mathrm{t}}$ and ${\delta}^{\mathrm{f}}$ are the STO and CFO after synchronization between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. The spatial characteristics are determined by azimuth/elevation angles of arrival (AoA) $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$ and departure (AoD) $\left\{\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right\}$. To simplify the notation, we further define
$\left\{\begin{array}[]{l}\varPhi_{k}^{\mathrm{r}}\triangleq\cos\theta_{k}^{\mathrm{r}},\,\varTheta_{k}^{\mathrm{r}}\triangleq\sin\theta_{k}^{\mathrm{r}}\cos\phi_{k}^{\mathrm{r}},\\ \varPhi_{k}^{\mathrm{t}}\triangleq\cos\theta_{k}^{\mathrm{t}},\,\varTheta_{k}^{\mathrm{t}}\triangleq\sin\theta_{k}^{\mathrm{t}}\cos\phi_{k}^{\mathrm{t}}.\\ \end{array}\right.$ (7)
@@ III-A Tensor Formulation of the Received Signal
By stacking the $N$ OFDM symbols and $M$ subcarriers, the received signal can be transformed to a three-order tensor $\bm{\mathcal{Y}}\in\mathbb{C}^{R\times N\times M}$ as follows
$\bm{\mathcal{Y}}=\sum_{k=1}^{K+1}{\alpha_{k}\mathbf{a}(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}})\circ\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})\circ\mathbf{c}(\tau_{k},{\delta}^{\mathrm{t}})}+\mathcal{N},$ (21)
where $\mathcal{N}$ is the noise tensor, $\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\in\mathbb{C}^{R\ \times 1}$, $\mathbf{b}\left(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}}\right)\in\mathbb{C}^{N\times 1}$, and $\mathbf{c}\left(\tau_{k},{\delta}^{\mathrm{t}}\right)\in\mathbb{C}^{M\times 1}$ can be respectively derived as follows
@@
$\bm{\mathcal{Y}}=\llbracket\mathbf{A},\mathbf{B},\mathbf{C}\rrbracket+\mathcal{N}.$ (25)
@@ III-B Decomposition-Based Factor Matrices Recovery
=$\overset{\left(a\right)}{=}\left(\mathbf{C}^{L_{1}}\odot\mathbf{B}\right)\left(\mathbf{C}^{L_{2}}\odot\mathbf{A}\right)^{\mathrm{T}}\in\mathbb{C}^{L_{1}N\times L_{2}R},$ (29)
@@
where $\mathbf{U}\in\mathbb{C}^{L_{1}N\times(K+1)}$, $\mathbf{\Sigma}\in\mathbb{C}^{(K+1)\times(K+1)}$, and $\mathbf{V}\in\mathbb{C}^{L_{2}R\times(K+1)}$. Note that $\mathbf{U}$ span the column spaces of $\mathbf{Y}_{S}$. Therefore, a nonsingular matrix $\mathbf{M}\in\mathbb{C}^{(K+1)\times(K+1)}$ exists such that [26]
@@ III-C Parameter Estimation
$\hat{\theta}_{k}^{\mathrm{r}}=\mathrm{arc}\cos\mathrm{(}\hat{\varPhi}_{k}^{\mathrm{r}}),\ \hat{\phi}_{k}^{\mathrm{r}}=\mathrm{arc}\cos\left(\frac{\hat{\varTheta}_{k}^{\mathrm{r}}}{\sin\mathrm{(}\hat{\theta}_{k}^{\mathrm{r}})}\right).$ (65)
@@
Algorithm 1 SVD-based Decoupling AoA Estimation Algorithm
@@ III-D Uniqueness Analysis
This subsection establishes uniqueness guarantees for the tensor decomposition (25) to ensure reliable recovery of factor matrices and subsequent parameter estimation. By extending the classical Kruskal’s uniqueness theorem [30] and incorporating structural constraints inherent to the factor matrices, we have the following relaxed uniqueness criterion [21, 31]:
@@
Consider a third-order tensor $\mathcal{X}\in\mathbb{C}^{I_{1}\times I_{2}\times I_{3}}$ with rank $K$ decomposed into factor matrices $\mathbf{A}^{(1)}\in\mathbb{C}^{I_{1}\times K}$, $\mathbf{A}^{(2)}\in\mathbb{C}^{I_{2}\times K}$ and $\mathbf{A}^{(3)}\in\mathbb{C}^{I_{3}\times K}$, where $\mathbf{A}^{(3)}$ is a Vandermonde matrix with generators $\{z_{k}\}_{k=1}^{K}$. Let $k(\mathbf{A})$ denote the Kruskal-rank of matrix $\mathbf{A}$. The decomposition achieves uniqueness if the following joint rank conditions hold:
$k\left(\mathbf{A}^{(L_{1},3)}\odot\mathbf{A}^{(2)}\right)=K,$ (66)
$k\left(\mathbf{A}^{(L_{2},3)}\odot\mathbf{A}^{(1)}\right)=K,$ (67)
where $\mathbf{A}^{(L_{1},3)}$ represents the first $L_{1}$ rows of $\mathbf{A}^{(3)}$, and $\mathbf{A}^{(L_{2},3)}$ denotes the first $L_{2}$ rows of $\mathbf{A}^{(2)}$. The uniqueness means that any possible combination of factor matrices $(\hat{\mathbf{A}}^{(1)},\mathrm{}\hat{\mathbf{A}}^{(2)},\hat{\mathbf{A}}^{(3)})$ for $\mathcal{X}$ satisfies
where $\mathbf{\Pi}$ is a $K\times K$ permutation matrix, and $\mathbf{\Delta}_{1}$, $\mathbf{\Delta}_{2}$, and $\mathbf{\Delta}_{3}$ with $\mathbf{\Delta}_{1}\mathbf{\Delta}_{2}\mathbf{\Delta}_{3}=\mathbf{I}_{K}$ are diagonal matrices. For generic scenarios, these conditions reduce to the simplified dimensionality constraint:
@@ III-E Algorithm Development
Algorithm 2 Tensor-Based Sensing Parameter Estimation Algorithm
@@ IV-A Basic Target Localization
where $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ is the horizontal orientation of rBS $n_{\mathrm{r}}$’s array, defined as the angle between the array normal and the global $x$-axis. Solving (73) for $\forall k$ yields basic position estimates $\{\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}\}_{k=1}^{K}$ related to tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. However, independent processing across the tBS-rBS pairs introduces data association ambiguity: due to inconsistent indexing across pairs, the estimated $\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}$ and $\hat{\mathbf{p}}_{n^{\prime}_{\mathrm{t}},k,n^{\prime}_{\mathrm{r}}}^{\mathrm{u}}$ may correspond to different targets when $n_{\mathrm{t}}\neq n^{\prime}_{\mathrm{t}}$ or $n_{\mathrm{r}}\neq n^{\prime}_{\mathrm{r}}$. In addition, estimation errors in $\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, $\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, and $\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$ lead to biases in position estimates. Although greedy algorithms can be applied for the target-to-estimate association [32], the erroneous position estimates cannot be effectively eliminated. This limitation often leads to incorrect data associations or significant global positioning biases, particularly in scenarios with a large number of targets or low signal-to-noise ratios (SNRs). In the following, an effective error elimination and data association algorithm is proposed based on the MST method.
@@ IV-B Error Elimination and Data Association
=$\mathcal{E}=\left\{\left(e_{1},e_{2}\right)|e_{1}\in V_{n_{\mathrm{t}},n_{\mathrm{r}}},\,e_{2}\in V\backslash V_{n_{\mathrm{t}},n_{\mathrm{r}}}\right\}.$ (81)
@@
[…]
@@
=$\mathcal{E}^{\prime}=\left\{\left(e_{1},e_{2}\right)\in\mathcal{E}|W_{\min}\left(e_{1}\right)>\varpi\right\},$ (83)
@@
Finally, the MST of the pruned graph $\tilde{\mathcal{G}}$ can be computed using the well-established algorithms such as Prim’s [33] or Kruskal’s [34]. Since the valid position estimates are clustered, the edges connecting estimates of the same target usually exhibit smaller weights and are preferentially retained in the MST. Therefore, the final data association is achieved by removing the $K-1$ longest edges from the MST. This operation partitions the MST into $K$ sub-graphs as shown in Fig. 3(c), each corresponding to a unique target.
@@ IV-E Algorithm Development
Algorithm 3 Cooperative Position and Velocity Estimation Algorithm
@@
9:Obtain sub-graphs by removing the $K-1$ longest edges from the MST;
@@
Complexity Analysis: The computational complexity of Algorithm 3 is dominated by the MST generation in Step 8 and solving Problem (87) in Step 12. The MST generation via Kruskal’s algorithm incurs a complexity of $𝑂\superscript\superscriptO\left(\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\log\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\right)$, where |\$\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\leqslant\left|\mathcal{E}\right|<\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)^{2}$. Given typical ISAC network scales, this simplifies to $O(\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)^{2}\log\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right))$. To solve Problem (87), the computational complexity of each quasi-Newton iteration is of $O\left(N_{\mathrm{t}}N_{\mathrm{r}}\right)$. With the averaging-based initialization $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}$ in (85), the required iteration number $G^{\prime}$ remains small (typically $G^{\prime}\leqslant 5$). Therefore, the total complexity of generating the fused position estimates $\left\{\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}\right\}_{k=1}^{K}$ is of $O\left(KG^{\prime}N_{\mathrm{t}}N_{\mathrm{r}}\right)$. To sum up, the total complexity of Algorithm 3 is given by $O((KN_{\mathrm{t}}N_{\mathrm{r}})^{2}\log\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)+KG^{\prime}N_{\mathrm{t}}N_{\mathrm{r}})$.
@@ V-A Simulation Setups
The simulation considers a cooperative ISAC system comprising 8 BSs uniformly distributed on a circle with a radius of 500 meters (m), as illustrated in Fig. 4. Each BS is positioned at a height of 30 m and equipped with a uniform planar array (UPA) consisting of 16 horizontal and 24 vertical antenna elements, driven by 64 radio frequency (RF) chains. The horizontal orientation $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ of each UPA is configured to point towards the center of the circle. To ensure the uniqueness of the CP decomposition, the elements of the precoding matrix $\mathbf{F}$ and the combining matrix $\mathbf{Q}$ are randomly selected from points on the complex unit circle [19].
@@
UAVs are modelled as sensing targets, generated uniformly at random within a circular area of radius 400 m and at heights ranging between 50 m and 300 m. The UAVs exhibit a maximum flight speed of 60 km/h and a minimum speed of 5 km/h, with a safety separation distance of 10 m enforced to prevent collisions. The radar cross-section (RCS) of each UAV is set to $\sigma_{\mathrm{RCS}}=0.1$ m2.
@@
where $d^{\mathrm{t}}$ and $d^{\mathrm{r}}$ are the distances in km from the target to the tBS and rBS, respectively, and $d^{\mathrm{LoS}}$ denotes the distance in km between the tBS and rBS.
@@ V-B Sensing Parameter Extraction Performance
Recall that all rBSs in the proposed sensing framework execute identical parameter extraction algorithms. To evaluate the parameter extraction performance of the proposed scheme, we first consider a bistatic sensing scenario with a single tBS-rBS pair deployed at (-500 m, 0) and (500 m, 0). The core challenge in estimating bistatic ranges $\left\{d_{k}\right\}$, azimuth/elevation angles $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$, and bistatic Doppler velocities $\left\{v_{k}\right\}$ is to recover the factor matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ by solving Problem (26). For performance benchmarking, we introduce a baseline scheme termed ALS. This approach solves Problem (26) using the ALS algorithm [19, 24], serving as a comparative reference to the proposed Algorithm 2.
@@
Fig. 5 compares the computational complexity of the proposed scheme and the ALS baseline, measured by CPU runtime versus the number of sensing targets $K$. The proposed algorithm demonstrates excellent computational efficiency, with its runtime showing almost negligible growth when the number of targets exceeds $K=3$. On the contrary, the ALS scheme requires higher computational resources in all scenarios, with its runtime increasing approximately exponentially as $K$ grows. Moreover, the proposed algorithm’s runtime at $K=6$ targets remains lower than ALS’s at $K=2$, highlighting its scalability advantage. The computational efficiency of this algorithm, which is unaffected by the growth in the number of targets, enables its practical deployment in real-time multi-target sensing applications, particularly in dense environments where system implementation is constrained by computational resources.
@@
Fig. 6 presents the parameter estimation RMSE performance for $K=\left\{2,3\right\}$ targets under varying transmit power at the tBS. It can be observed that both schemes show monotonic estimation accuracy improvement with increased tBS transmit power, resulting from the increment of SNR. For all sensing parameters, the proposed algorithm demonstrates significant advantages in accuracy, particularly in bistatic range extraction, where the advantage reaches one to two orders of magnitude, achieving millimetre-level precision under sufficient transmit power. Conversely, the ALS scheme suffers from slow and unstable convergence, with limited estimation accuracy. Based on these parameter estimates, Fig. 7 evaluates the RMSE of the basic target position estimation derived from (73). Thanks to the high-precision parameter extraction ability of Algorithm 2, the proposed scheme achieves a localization accuracy approximately one order of magnitude higher than the ALS scheme, with a decimetre-level positioning accuracy when the transmit power is higher than 55 dBm.

### v2

@@ II-C Sensing Channel Model
where $\alpha_{k}$ represents the composite channel gain of the $k$th path, while ${\delta}^{\mathrm{t}}$ and ${\delta}^{\mathrm{f}}$ are the STO and CFO after synchronization between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. The channel gain is influenced by factors such as the path length $d_{k}$ and radar cross-section (RCS) $\sigma_{\mathrm{RCS}}$ of targets, making direct estimation challenging. The spatial characteristics are determined by elevation/azimuth AoA $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$ and AoD $\left\{\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right\}$. To simplify the notation, we further define
[…]
@@ III-A Tensor Formulation of the Received Signal
By stacking the $N$ OFDM symbols and $M$ subcarriers, the received signal can be transformed to a three-order tensor $\bm{\mathcal{Y}}\in\mathbb{C}^{R\times N\times M}$ as follows [25]
$\bm{\mathcal{Y}}=\sum_{k=1}^{K+1}{\alpha_{k}\mathbf{a}(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}})\circ\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})\circ\mathbf{c}(\tau_{k},{\delta}^{\mathrm{t}})}+\bm{\mathcal{N}},$ (18)
where $\bm{\mathcal{N}}$ is the noise tensor, $\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\in\mathbb{C}^{R\times 1}$, $\mathbf{b}\left(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}}\right)\in\mathbb{C}^{N\times 1}$, and $\mathbf{c}\left(\tau_{k},{\delta}^{\mathrm{t}}\right)\in\mathbb{C}^{M\times 1}$ can be respectively derived as follows
@@
$\bm{\mathcal{Y}}=\llbracket\mathbf{A},\mathbf{B},\mathbf{C}\rrbracket+\bm{\mathcal{N}}.$ (22)
@@ III-B Decomposition-Based Factor Matrices Recovery
$\overset{\left(a\right)}{=}\left(\mathbf{C}^{L_{1}}\odot\mathbf{B}\right)\left(\mathbf{C}^{L_{2}}\odot\mathbf{A}\right)^{\mathrm{T}}\in\mathbb{C}^{L_{1}N\times L_{2}R},$ (26)
@@
where $\mathbf{U}\in\mathbb{C}^{L_{1}N\times(K+1)}$, $\mathbf{\Sigma}\in\mathbb{C}^{(K+1)\times(K+1)}$, and $\mathbf{V}\in\mathbb{C}^{L_{2}R\times(K+1)}$. Note that $\mathbf{U}$ spans the column spaces of $\mathbf{Y}_{S}$. Therefore, a nonsingular matrix $\mathbf{M}\in\mathbb{C}^{(K+1)\times(K+1)}$ exists such that [28]
@@ III-C Parameter Estimation
$\hat{\theta}_{k}^{\mathrm{r}}=\mathrm{arc}\cos(\hat{\varPhi}_{k}^{\mathrm{r}}),\mathrm{}\hat{\phi}_{k}^{\mathrm{r}}=\mathrm{arc}\cos\left(\frac{\hat{\varTheta}_{k}^{\mathrm{r}}}{\sin(\hat{\theta}_{k}^{\mathrm{r}})}\right).$ (62)
@@
Algorithm 1 SVD-based Decoupling AoA Estimation
@@ III-D Uniqueness Analysis
This subsection establishes uniqueness guarantees for the tensor decomposition (22) to ensure reliable recovery of factor matrices. Consider a third-order tensor $\mathcal{X}\in\mathbb{C}^{I_{1}\times I_{2}\times I_{3}}$ with rank $K$ decomposed into factor matrices $\mathbf{A}^{(1)}\in\mathbb{C}^{I_{1}\times K}$, $\mathbf{A}^{(2)}\in\mathbb{C}^{I_{2}\times K}$ and $\mathbf{A}^{(3)}\in\mathbb{C}^{I_{3}\times K}$, where $\mathbf{A}^{(3)}$ is a Vandermonde matrix with generators $\{z_{k}\}_{k=1}^{K}$. The uniqueness means that any possible combination of factor matrices $(\hat{\mathbf{A}}^{(1)},\mathrm{}\hat{\mathbf{A}}^{(2)},\hat{\mathbf{A}}^{(3)})$ for $\mathcal{X}$ satisfies
where $\mathbf{\Pi}$ is a $K\times K$ permutation matrix, and $\mathbf{\Delta}_{1}$, $\mathbf{\Delta}_{2}$, and $\mathbf{\Delta}_{3}$ with $\mathbf{\Delta}_{1}\mathbf{\Delta}_{2}\mathbf{\Delta}_{3}=\mathbf{I}_{K}$ are diagonal matrices. By extending the classical Kruskal’s uniqueness theorem [32] and incorporating structural constraints inherent to the factor matrices, the following relaxed uniqueness criterion has been proved in [33, Theorem III.3]:
@@
Let $k(\mathbf{A})$ denote the Kruskal-rank of matrix $\mathbf{A}$. The tensor decomposition of $\mathcal{X}$ is unique if the following joint rank conditions hold:
$k(\mathbf{A}^{(L_{1},3)}\odot\mathbf{A}^{(2)})=K,$ (64)
$k(\mathbf{A}^{(L_{2},3)}\odot\mathbf{A}^{(1)})=K,$ (65)
where $\mathbf{A}^{(L_{1},3)}$ represents the first $L_{1}$ rows of $\mathbf{A}^{(3)}$, and $\mathbf{A}^{(L_{2},3)}$ denotes the first $L_{2}$ rows of $\mathbf{A}^{(2)}$. These conditions are generically true if
@@ III-E Algorithm Development
Algorithm 2 Tensor-Based Sensing Parameter Estimation
@@ IV-A Basic Target Localization
where $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ is the horizontal orientation of rBS $n_{\mathrm{r}}$’s array, defined as the angle between the array normal and the global $x$-axis. Solving (70) for $\forall k$ yields basic position estimates $\{\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}\}_{k=1}^{K}$ related to tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. However, independent processing across the tBS-rBS pairs introduces data association ambiguity: due to inconsistent indexing across pairs, the estimated $\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}$ and $\hat{\mathbf{p}}_{n^{\prime}_{\mathrm{t}},k,n^{\prime}_{\mathrm{r}}}^{\mathrm{u}}$ may correspond to different targets when $n_{\mathrm{t}}\neq n^{\prime}_{\mathrm{t}}$ or $n_{\mathrm{r}}\neq n^{\prime}_{\mathrm{r}}$. In addition, estimation errors in $\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, $\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, and $\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$ lead to biases in position estimates. Although greedy algorithms can be applied for the target-to-estimate association [34], the erroneous position estimates cannot be effectively eliminated. This limitation often leads to incorrect data associations or significant global positioning biases, particularly in scenarios with a large number of targets or low SNRs. In the following, an effective error elimination and data association algorithm is proposed based on the MST method.
@@ IV-B Error Elimination and Data Association
$\mathcal{E}=\left\{\left(e_{1},e_{2}\right)|e_{1}\in V_{n_{\mathrm{t}},n_{\mathrm{r}}},\,e_{2}\in V\backslash V_{n_{\mathrm{t}},n_{\mathrm{r}}}\right\}.$ (78)
@@
where $\xi^{\mathrm{t}}\left(e\right)=\mathrm{mod}\left(\mathrm{mod}\left(e-1,K\right),N_{\mathrm{t}}\right)+1$, $\xi^{\mathrm{u}}\left(e\right)=\lfloor\frac{e_{1}-1}{K}\rfloor+1$ and $\xi^{\mathrm{r}}\left(e\right)=\lfloor\frac{\mathrm{mod}\left(e_{1}-1,K\right)}{N_{\mathrm{t}}}\rfloor+1$ decodes the tBS-target-rBS indices from vertex $e$, respectively, with $\mathrm{mod}$ and $\lfloor\cdot\rfloor$ denoting the modulus and floor operations, respectively.
@@
$\mathcal{E}^{\prime}=\left\{\left(e_{1},e_{2}\right)\in\mathcal{E}|W_{\min}\left(e_{1}\right)>\varpi\right\},$ (80)
@@
Finally, the MST of the pruned graph $\tilde{\mathcal{G}}$ can be computed using the well-established algorithms such as Prim’s [35] or Kruskal’s [36]. Since the valid position estimates are clustered, the edges connecting estimates of the same target usually exhibit smaller weights and are preferentially retained in the MST. Therefore, the final data association is achieved by removing the K1K-1 longest edges from the MST. This operation partitions the MST into $K$ sub-graphs as shown in Fig. 4(c), each corresponding to a unique target.
@@ IV-E Algorithm Development
Algorithm 3 Cooperative Position and Velocity Estimation
@@
9:Obtain sub-graphs by removing the K1K-1 longest edges from the MST;
@@
[…]
@@ V-A Simulation Setups
The simulation considers a cooperative ISAC system comprising 8 BSs uniformly distributed on a circle with a radius of 500 meters (m), as illustrated in Fig. 5. Each BS is positioned at a height of 30 m and equipped with a uniform planar array (UPA) consisting of 16 horizontal and 24 vertical antenna elements, driven by 64 radio frequency (RF) chains. The horizontal orientation $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ of each UPA is configured to point towards the center of the circle. The precoding matrices $\mathbf{F}$ and the combining matrices $\mathbf{Q}$ for all BSs are constructed following the exemplary design in Section II-B.
@@
UAVs are modelled as sensing targets, generated uniformly at random within a circular area of radius 400 m and at heights ranging between 50 m and 300 m. The UAVs exhibit a maximum flight speed of 60 km/h and a minimum speed of 5 km/h, with a safety separation distance of 10 m enforced to prevent collisions. The RCS of each UAV is set to $\sigma_{\mathrm{RCS}}=0.01$ m2.
@@
where $d^{\mathrm{t}}$ and $d^{\mathrm{r}}$ are the distances in km from the target to the tBS and rBS, respectively, and $d^{\mathrm{LoS}}$ denotes the distance in km between the tBS and rBS. The STO and CFO between each tBS-rBS pair are set to $10^{-8}$ s and $0.01\Delta f$, respectively.
@@ V-B Sensing Parameter Extraction Performance
Recall that all rBSs in the proposed sensing framework execute identical parameter extraction algorithms. To evaluate the parameter extraction performance of the proposed scheme, we first consider a bistatic sensing scenario with a single tBS-rBS pair deployed at (-500 m, 0) and (500 m, 0). The core challenge in estimating bistatic ranges $\left\{d_{k}\right\}$, azimuth/elevation angles $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$, and bistatic Doppler velocities $\left\{v_{k}\right\}$ is to recover the factor matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ by solving Problem (23). For performance benchmarking, we introduce a baseline scheme termed ALS [^3] Although other parameter extraction algorithms such as 2D-MUSIC or 2D-FFT exist, they estimate each set of parameters separately and cannot directly match parameters belonging to the same target in multi-target scenarios. ALS is adopted as it is the only conventional method capable of jointly estimating all parameters with correct pairing.. This approach solves Problem (23) using the ALS algorithm [19, 26], serving as a comparative reference to the proposed Algorithm 2.
@@
Fig. 6 compares the computational complexity of the proposed scheme and the ALS baseline, measured by CPU runtime versus the number of sensing targets $K$. The proposed algorithm demonstrates excellent computational efficiency, with its runtime showing almost negligible growth when the number of targets exceeds $K=3$. On the contrary, the ALS scheme requires higher computational resources in all scenarios, with its runtime increasing approximately exponentially as $K$ grows. Moreover, the computational efficiency of the proposed algorithm is not significatly affected by the growth in the number of targets. This indicates the scalability advantage of the proposed algorithm in real-time multi-target sensing applications, particularly in dense environments where system implementation is constrained by computational resources.
@@
Fig. 7 presents the parameter estimation RMSE performance for $K=\left\{2,4\right\}$ targets under varying transmit power at the tBS. In each subfigure, a horizontal line is included to indicate the theoretical resolution limit for the corresponding sensing parameter, which is determined by the system configuration. It can be observed that both schemes exhibit monotonic improvements in estimation accuracy with increasing tBS transmit power, resulting from the increment of SNR. The proposed algorithm demonstrates significant advantages in estimation accuracy across all parameters, particularly in bistatic range estimation, where it outperforms the baseline by one to two orders of magnitude, achieving millimeter-level precision under sufficient transmit power. Conversely, the ALS scheme shows limited estimation accuracy due to slow and unstable convergence. Interestingly, thanks to the joint extraction of time delays, Doppler shifts, and angles from the received signal tensor $\bm{\mathcal{Y}}$, the proposed scheme achieves an RMSE far below the theoretical resolution limit, especially in Doppler velocity estimation. It verifies that the proposed algorithm achieves considerable sensing accuracy with only a small number of OFDM symbols ($N=7$), thereby minimizing the impact on the system’s communication capacity.
Based on these parameter estimates, Fig. 8 evaluates the RMSE of the basic target position estimation derived from (70). Thanks to the high-precision parameter extraction ability of Algorithm 2, the proposed scheme achieves a localization accuracy approximately one order of magnitude higher than the ALS scheme, with a decimetre-level positioning accuracy when the transmit power reaches 55 dBm.